OpenStudy (anonymous):

What number must you add to complete the square? x2 + 12x = 33 helP!

OpenStudy (raden):

added by (1/2 * 12)^2 = 6^2 = 36

OpenStudy (amistre64):

http://openstudy.com/study#/updates/51a6698be4b0ffa8e49a2edf just in case ....

OpenStudy (anonymous):

omg can i get answer plz

OpenStudy (amistre64):

no

OpenStudy (anonymous):

idk man i forgot 9th grade stuff

OpenStudy (anonymous):

I've never learn this before

OpenStudy (amistre64):

me too, had to relearn it all in my 40s

OpenStudy (amistre64):

the setup follows the rundown in the link ... we are essentially try to find a value of n^2, such that 2n = 12 in this case

OpenStudy (amistre64):

2, times what, equals 12?

OpenStudy (anonymous):

4

OpenStudy (amistre64):

almost, 2*4 = 8 ... lets assume 2*6 = 12

OpenStudy (amistre64):

since n=6, and we want n^2; what is the value of 6^2 ?

OpenStudy (anonymous):

i do not have any calculator on me

OpenStudy (anonymous):

u divide or something

OpenStudy (amistre64):

you might need to go back and brush up on your basic times tables then ... or you have forgotten what ^2 means.

OpenStudy (amistre64):

n^2 is notation for n times n 6^2 is notation for 6 times 6

OpenStudy (anonymous):

36

OpenStudy (amistre64):

correct :) so, in order to "complete the square" we need to add 36 to both sides of your setup

OpenStudy (anonymous):

so 36 is the complete square

OpenStudy (amistre64):

no, the value "36" completes the square for us x2 + 12x = 33 x2 + 12x +36 = 33 +36 (x-6)^2 = 69

OpenStudy (anonymous):

?

OpenStudy (anonymous):

oh k

OpenStudy (anonymous):

69

OpenStudy (anonymous):

man thats a lot of work

OpenStudy (amistre64):

it is, which is why computers are sooo great these days :)

OpenStudy (amistre64):

you eventually just get into the habit of squaring half of the coefficient of x

OpenStudy (amistre64):

What number must you add to complete the square? x2 + 12x = 33 ^^^ 12/2 = 6 6^2 = 36

OpenStudy (anonymous):

so thats the answer 36

OpenStudy (amistre64):

i hope so .... unless weve been working on a different problem :/

OpenStudy (anonymous):

anyway thanks man i just took a quiz on this and i failed it

OpenStudy (amistre64):

it takes a bit of practice to get down pat, and you really have no reason to come across it when working out there in the real world.

OpenStudy (anonymous):

i hate doing this